OpenStudy (anonymous):

How can we explain using our knowledge of components, in words, the fact that the speed with which a projectile is launched is the same as the speed just before it strikes the ground? Ignore air resistance and assume the launch takes place on flat ground.

OpenStudy (anonymous):

the horizontal component remains same , as no acceleration is present in that direction . also , it is known that is u thrw a body up with a certain vel , it comes back with same vel...so the vertical component at the moment of striking the ground bcoms equal to that when it was thrown...so as both components are same the vel is same....note that the direction of vertical component changes so does the direction of the velocity

OpenStudy (anonymous):

Mathematically : Horizontally we have \[u \cos \theta=constant\] Vertically we have: \[v ^{2}=u ^{2}+2.a.S\] or,\[v^2=u^2Sin^2\theta + 2.(-g).S\] or,S=0 as vertical displ is zero so the components are same!!